Suppose you observe a star orbiting the galactic center at a speed of 1000 km/s in a circular orbit with a radius of 16 light-days.

Calculate the mass of the object that the star is orbiting.
Answer #1

We know
Speed of light = 300,000km/s
Seconds in a day = 86400
Given Seconds in 16 days = 16*86400=1382400

Radius = 300,000 x 1382400 = 4.1472*10^11 km.

We need m/s and metres though, so multiply what we need by 1000.

M = (1000x1000)^2 x (5.184x10^11 x 1000) / 6.67x10^-11

M = 6.217x10^36kg

Hope u like it